OpenStudy (anonymous):

simplify 1/(1-x^2)^-1/2

OpenStudy (anonymous):

remember that \[\frac{1}{a^{-n}} = a^n\] thus, u can simplify it to \[\sqrt{1-x^2}\]

OpenStudy (anonymous):

thank you.